# Natural Language Processing in Patents: A Survey

### v2

@@ A.2 Patent Retrieval
A.3 Figure Caption Generation
Researchers studied generating brief descriptions of the patent figures. Aubakirova et al. (2023) introduced the first large-scale patent figure dataset, comprising more than 30,000 figure–caption pairs. The authors fine-tuned large vision language models to generate captions as baselines on this novel dataset and call for future research on the improvement of existing caption generation models. As contextual information, the quality of such brief figure descriptions can vary significantly. In quite many cases, the figure captions read Figure n depicts another embodiment of the invention or Figure m depicts perspective xy of an embodiment of the invention to fulfill the minimum formalities (e.g., per European Patent Regulations Rule 42(d), PCT Regulations Rule 5.1(a)(iv), US 37 CFR 1.74) and let the subsequent detailed description of the figures and embodiments provide more content.
@@ B.3 Text Generation
Recently, Lee (2023) proposed a novel keystroke-based evaluation method for generative patent models. This method measured the number of keystrokes that the model could save by providing model predictions in an auto-complete function. The experimental results demonstrated that the performance of larger models was not better by using this evaluation metric. However, the experiments only focused on patent claims generation and the effectiveness of this metric on other generative tasks was unknown.

### v3

@@ A.2 Patent Retrieval
A.3 Figure Caption/Description Generation
Researchers studied generating brief descriptions of the patent figures. Aubakirova et al. (2023) introduced the first large-scale patent figure dataset, comprising more than 30,000 figure–caption pairs. The authors fine-tuned large vision language models to generate captions as baselines on this novel dataset and call for future research on the improvement of existing caption generation models. As contextual information, the quality of such brief figure descriptions can vary significantly. In quite many cases, the figure captions read Figure n depicts another embodiment of the invention or Figure m depicts perspective xy of an embodiment of the invention to fulfill the minimum formalities (e.g., per European Patent Regulations Rule 42(d), PCT Regulations Rule 5.1(a)(iv), US 37 CFR 1.74) and let the subsequent detailed description of the figures and embodiments provide more content. Similarly, Shukla et al. (2025) presented PatentDesc, a novel large-scale dataset that comprises approximately 355K patent figures paired with both brief and detailed textual descriptions. They trained a vision encoder specifically designed for patent figures and demonstrated that it can significantly enhance the generation performance. Furthermore, Shomee et al. (2024) introduced IMPACT, a multimodal patent dataset that contains 3.61 million design patent figures paired with detailed captions and rich metadata. They used the dataset to build benchmarks for two novel computer vision tasks—3D image construction and visual question answering (VQA).
@@ B.3 Text Generation
Recently, Lee (2023) proposed a novel keystroke-based evaluation method for generative patent models. This method measured the number of keystrokes that the model could save by providing model predictions in an auto-complete function. The experimental results demonstrated that the larger models could not improve this specific evaluation metric. However, the experiments only focused on patent claims generation and the effectiveness of this metric on other generative tasks was unknown.
A recent work proposed PatentEval as a human evaluation framework to evaluate machine-generated patent texts (Zuo et al., 2024). The framework categorizes errors into different types, such as grammatical errors, irrelevant content, incomplete coverage, and clarity issues. The authors manually assessed 400 granted patents and indicated that while LLMs demonstrate potential for patent drafting, they still exhibit limitations.